The length of a rectangle is 8 cm less than twice the width. The area of the rectangle is 90 cm2. Find the length and width of the rectangle.
length______cm
width_______cm

Answer by jorel1380(3719)

n(2n-8)=90
2n2-8n-90=0
n2-4n-45=0
(n-9)(n+5)=9
n=9 or -5
Throwing out the negative answer, we get the length to be 10 cm., and the width to be 9 cm..